OpenStudy (anonymous):

How could I solve (x+5)/(x^2+x) = 1/(x^2+x) - (x-6)/(x+1)

OpenStudy (lgbasallote):

multiply everything by an LCM which in this case is (x^2 +x)(x+1) so... (x+5)(x+1) = x+1 - (x-6)(x^2+x) x^2 + 6x + 5 = x + 1 - (x^3 -5x^2-6x) x^2 + 6x + 5 = x + 1 - x^3 +5x^2 +6x x^2 + 6x + 5 - x - 1 + x^3 - 5x^2 - 6x = 0 x^3 -4x^2 - x + 4 = 0 are you looking for x?

OpenStudy (anonymous):

the problem just said solve :P